# Wi-Fi Radar via Over-the-Air Referencing: Bridging Wi-Fi Sensing and Bistatic Radar

### v1

@@ Abstract
Wi-Fi sensing has attracted significant attention for human sensing and related applications. However, unsynchronized transmitters and receivers fundamentally preclude phase-coherent radar-like delay–Doppler analysis. By exploiting the line-of-sight (LoS) path, i.e., the earliest-arriving direct path, as an over-the-air (OTA) reference for delay and phase, we propose an OTA LoS-path referencing scheme, termed LoSRef, that enables delay calibration and phase alignment in unsynchronized Wi-Fi systems. Unlike conventional Wi-Fi bistatic radar systems that rely on wired reference signals or dedicated reference antennas, the proposed LoSRef-based framework bridges the long-standing gap between conventional Wi-Fi sensing and Wi-Fi radar, enabling phase-coherent bistatic radar-like operation in a drop-in Wi-Fi sensing configuration. Through human gait and respiration experiments in indoor environments, we demonstrate that phase-coherent channel impulse responses and corresponding delay–Doppler responses are obtained using only commodity Wi-Fi devices. This enables physically interpretable human motion sensing, including gait-induced range variation and respiration-induced sub-wavelength displacement, as well as the extraction of target-induced dynamics up to 20 dB weaker than dominant static multipath components.
@@ I Introduction
As a representative example of integrated sensing and communication (ISAC) that leverages existing wireless infrastructure [1], device-free Wi-Fi sensing has attracted significant attention in recent years [2]. By exploiting ubiquitous Wi-Fi transmissions, numerous studies have demonstrated the feasibility of monitoring human respiration, gait, and daily activities without requiring the target to carry any dedicated device [3, 4]. These capabilities have enabled a broad range of applications in healthcare, smart living environments, and ambient intelligence, while benefiting from the low deployment cost and widespread availability of commercial Wi-Fi networks.
Early Wi-Fi sensing studies primarily relied on coarse metrics such as received signal strength indicator (RSSI) [5, 6]. A major turning point was the release of the Linux 802.11n channel state information (CSI) tool [7], which enabled researchers to access CSI from commercial off-the-shelf (COTS) IEEE 802.11n devices. In this context, CSI refers to the complex-valued channel frequency response (CFR) measured across orthogonal frequency division multiplexing (OFDM) subcarriers. Throughout this paper, the terms CSI and CFR are used interchangeably: CSI is used when referring to channel measurement and acquisition, while CFR denotes the frequency-domain representation of the channel and its Fourier-transform relationship to the channel impulse response (CIR), which represents the channel in the delay domain.
This CSI acquisition capability has triggered extensive research on CSI-based Wi-Fi sensing, including motion-related sensing based on temporal variations of the wireless channel. More recently, Wi-Fi sensing has matured beyond ad-hoc research prototypes, culminating in the standardization of IEEE 802.11bf, which defines channel measurement procedures to explicitly support a wide range of sensing use cases, including range, motion, and gesture detection, within Wi-Fi networks [8].
@@
Alongside the aforementioned Wi-Fi sensing approaches, there also exists a line of research that exploits COTS Wi-Fi transmissions for radar applications [9, 10]. For clarity, we refer to methods operating without explicit reference signals as Wi-Fi sensing, and those employing explicit reference signals for phase-coherent analysis as Wi-Fi radar. Early Wi-Fi radar studies typically relied on measurement-grade Rxs and external reference acquisition mechanisms, such as reference antennas or wired signal taps, as illustrated in Fig. 1(a) and (b), to enable coherent delay or Doppler estimation [9, 11, 10, 12, 13]. More recent works have demonstrated that reference signals can also be exploited in fully COTS-based systems to stabilize CSI measurements, i.e., the phase of the underlying CFR [14, 15, 16, 17]. These reference-aided architectures are fundamentally consistent with bistatic radar principles and clearly demonstrate the potential of Wi-Fi signals for radar-like sensing.
However, due to their reliance on explicit reference acquisition mechanisms, most prior Wi-Fi radar systems cannot operate in typical drop-in Wi-Fi sensing scenario as they require cable routing or dedicated reference antenna placement, thereby imposing additional constraints on deployment and installation beyond COTS Wi-Fi devices.
These observations reveal a fundamental gap between existing Wi-Fi sensing and Wi-Fi radar approaches. While reference-aided Wi-Fi radar enables physically meaningful delay and Doppler analysis, its reliance on auxiliary reference hardware limits the practicality of drop-in Wi-Fi deployments. Conversely, Wi-Fi sensing approaches that rely, for example, on CFR magnitude alone offer ease of deployment but lack direct access to physically interpretable propagation parameters.
This raises a key question: Can bistatic Wi-Fi radar functionality be realized using only unmodified COTS Wi-Fi devices, without wired references or dedicated reference antennas, while still enabling physically meaningful delay–Doppler estimation?
@@
  • Unlike most existing Wi-Fi sensing approaches that operate directly on the CFR, a channel impulse response (CIR) is constructed using CFR obtained from COTS IEEE 802.11ax devices with $160\text{\,}\mathrm{MHz}$ bandwidth [18]. This is enabled by carefully designed frequency-domain preprocessing of the raw CFR, and provides a reliable basis for subsequent phase-coherent delay-domain processing.
  • To the best of our knowledge, this is the first work that realizes reference-aided bistatic Wi-Fi radar using only unmodified COTS Wi-Fi devices, without relying on wired references or dedicated reference antennas. To this end, we propose an over-the-air (OTA) line-of-sight (LoS)-path referencing scheme (LoSRef) that exploits the temporal stability of the transmitter–receiver (Tx–Rx) LoS path to provide a common delay and phase reference in practical Wi-Fi systems, as illustrated in Fig. 1(c), thereby enabling phase-coherent CIR acquisition.
  • The proposed LoSRef-based framework requires only a single antenna port at the Rx, whereas conventional reference-aided bistatic Wi-Fi radar systems typically rely on multi-port Rxs or dedicated reference channels. This further simplifies hardware requirements and facilitates drop-in deployment with COTS Wi-Fi devices.
@@
This paper is organized as follows. Section II reviews related work on Doppler-based Wi-Fi sensing and reference-aided Wi-Fi radar. Section III discusses the reference ambiguity inherent in unsynchronized Wi-Fi systems. Section IV presents the proposed phase-coherent CIR construction framework, including delay and phase referencing based on a stable LoS-path reference. Section V demonstrates bistatic range estimation through human gait and respiration experiments. Section VI presents delay–Doppler response estimation results for moving targets. Section VII provides additional discussion, followed by Section VIII, which concludes the paper.
@@ II-A Doppler-Based Wi-Fi Sensing
TABLE I: Positioning of This Work with Respect to Existing Studies
Study Tx Rx Reference (Fig. 1) Standard
Typical Wi-Fi sensing [23, 3, 24] COTS COTS None IEEE 802.11n
[10, 12, 13] COTS Instrument Rx (a) Wired reference IEEE 802.11g
[9, 11] COTS Instrument Rx (b) Reference antenna IEEE 802.11g
[14, 16] COTS COTS (a) Wired reference IEEE 802.11n
[15] COTS COTS (b) Reference antenna IEEE 802.11n
[17] COTS COTS (b) Reference antenna IEEE 802.11ax
@@ II-B Reference-Aided Wi-Fi Radar
Table I summarizes the positioning of this work with respect to representative Wi-Fi sensing and reference-aided Wi-Fi radar studies, highlighting differences in Tx and Rx types, reference scheme, and supported standards. In [9, 11], a bistatic radar architecture is adopted, where a COTS Wi-Fi Tx and a measurement-grade Rx are used. The Rx employs separate reference and surveillance antennas, as illustrated in Fig. 1(b), with the antenna configuration designed such that target-induced variations are primarily captured in the Tx–surveillance path. The target-induced Doppler frequency is then extracted by correlating the reference and surveillance signals.
@@
In contrast to existing reference-aided Wi-Fi radar systems, which rely on wired references or dedicated reference antennas, the LoSRef exploits the OTA Tx–Rx LoS path. As a result, phase-coherent bistatic radar operation is achieved using only a single-antenna, single-port Rx, without requiring multi-port Rx architectures.
III CIR Representation and Reference Ambiguity in Unsynchronized Wi-Fi Systems
This section briefly reviews the CIR representation and highlights the reference ambiguity inherent in unsynchronized Wi-Fi systems, which fundamentally prevents phase-coherent analysis and physically meaningful time-domain processing.
@@ III-A CIR Representation and Doubly Selective Channels
III-B Reference Ambiguity in Unsynchronized Wi-Fi Systems
In Wi-Fi systems, where the Tx and Rx are not synchronized, neither a common timing reference nor an absolute phase reference is available between them. As a result, the CIR observed at the Rx does not preserve the absolute delay reference to the transmission timing, nor does it retain a phase-coherent representation across successive observations. Accordingly, the observed CIR $h_{\mathrm{raw}}(\tau,t)$ is related to the underlying physical CIR $h(\tau,t)$ through an unknown delay offset and an unknown phase offset, and is given by
@@ III-B Reference Ambiguity in Unsynchronized Wi-Fi Systems
where $\alpha_{t}$ and $\beta_{t}$ denote the delay offset and phase offset, respectively, which may vary across observation times $t$ [21, 22]. These offsets originate from free-running clocks and carrier phase differences between the Tx and Rx, and prevent direct use of absolute delay and phase information, thereby precluding phase-coherent analysis of the doubly selective channel in unsynchronized Wi-Fi systems. Note that the delay offset $\alpha_{t}$ of the devices used in this study has been experimentally characterized in [17].
@@ IV-B LoSRef-Based Delay Calibration and Phase Alignment
Figure 3: LoSRef: OTA LoS-path referencing for delay calibration and phase alignment. The dominant LoS-path component is used as an OTA reference to compensate unknown offsets caused by unsynchronized packet transmissions.
@@ IV-C Clutter Removal via LoSRef-based CIR Alignment
Note that, in conventional Wi-Fi sensing systems, direct processing of complex-valued CIRs is difficult due to the lack of delay synchronization between the Tx and Rx, which prevents meaningful temporal averaging. For example, Widar [24] extracts motion-related components by applying PCA to the CFR magnitude, followed by time–frequency spectral analysis. By contrast, the proposed LoSRef-based approach aligns CIRs on a common delay axis using a stable LoS-path reference, enabling phase-coherent temporal averaging and delay-domain clutter removal, which may be advantageous for slowly varying motions.
@@ VIII Conclusion
Through representative human gait and respiration experiments, we confirmed that phase-coherent bistatic radar functionality can be realized within a conventional device-free Wi-Fi sensing configuration. Specifically, the proposed framework enables physically meaningful delay and Doppler estimates, including gait direction discrimination and respiration-induced sub-wavelength displacement measurements.

### v2

@@ Abstract
Wi-Fi channel state information (CSI), which is originally acquired for communication purposes, has recently been reused for sensing and radar-like functionalities. However, in practical Wi-Fi systems with independent clocks at the transmitter and receiver, the lack of a common delay and phase reference fundamentally precludes phase-coherent radar-like delay–Doppler analysis. By exploiting the line-of-sight (LoS) path component, i.e., the earliest-arriving direct path, as an over-the-air (OTA) reference for delay and phase, we propose an OTA LoS-path referencing scheme, termed LoSRef, that enables delay calibration and phase alignment under this practical constraint. Unlike conventional Wi-Fi bistatic radar systems that rely on wired reference signals or dedicated reference antennas, the proposed LoSRef-based framework enables phase-coherent bistatic radar-like operation that can be integrated into typically deployed Wi-Fi systems. Through human gait and respiration experiments in indoor environments, we demonstrate that phase-coherent channel impulse responses and corresponding delay–Doppler responses can be obtained using only commodity Wi-Fi devices. This enables physically interpretable human motion sensing, including gait-induced range variation and respiration-induced sub-wavelength displacement, as well as the extraction of target-induced dynamics up to 20 dB weaker than dominant static multipath components.
@@ I Introduction
As a key application of integrated sensing and communications (ISAC), which aims to augment wireless communication systems with sensing capabilities by leveraging existing wireless infrastructure [1], wireless LAN systems with device-free sensing capability have attracted significant attention in recent years [2, 3, 4]. By exploiting ubiquitous Wi-Fi transmissions, numerous studies have demonstrated the feasibility of monitoring human respiration, gait, and daily activities without requiring the target to carry any dedicated device. These capabilities have enabled a broad range of applications in healthcare, smart living environments, and ambient intelligence, while benefiting from the low deployment cost and widespread availability of commercial Wi-Fi networks.
Early Wi-Fi sensing studies primarily relied on coarse metrics such as received signal strength indicator (RSSI) [5, 6]. A major turning point was the release of the Linux 802.11n channel state information (CSI) tool [7], which enabled researchers to access CSI from commercial off-the-shelf (COTS) IEEE 802.11n devices. In this context, CSI refers to the complex-valued channel frequency response (CFR) measured across orthogonal frequency division multiplexing (OFDM) subcarriers. From a communication-system perspective, CSI is obtained for communication purposes such as equalization, precoding, and link adaptation, rather than being designed specifically for sensing purposes. Throughout this paper, the terms CSI and CFR are used interchangeably: CSI is used when referring to channel measurement and acquisition, while CFR denotes the frequency-domain representation of the channel and its Fourier-transform relationship to the channel impulse response (CIR), which represents the channel in the delay domain.
This CSI acquisition capability has triggered extensive research on CSI-based Wi-Fi sensing, including motion-related sensing based on temporal variations of the wireless channel, as summarized in representative surveys such as [2, 3, 4]. More recently, Wi-Fi sensing has matured beyond ad-hoc research prototypes, culminating in the standardization of IEEE 802.11bf [8], which defines channel measurement procedures to explicitly support a wide range of sensing use cases, including range, motion, and gesture detection, within Wi-Fi networks.
@@
Alongside the aforementioned Wi-Fi sensing approaches, there also exists a line of research that exploits COTS Wi-Fi transmissions for radar applications [9, 10]. For clarity, we refer to methods operating without explicit reference signals as Wi-Fi sensing, and those employing explicit reference signals for phase-coherent analysis as Wi-Fi radar. Early Wi-Fi radar studies typically relied on measurement-grade receivers and external reference acquisition mechanisms, such as reference antennas or wired signal taps, as illustrated in Fig. 1(a) and (b), to enable coherent delay or Doppler estimation [9, 11, 10, 12, 13]. More recent works have demonstrated that reference signals can also be exploited in fully COTS-based systems to stabilize CSI measurements, i.e., the phase of the underlying CFR [14, 15, 16, 17]. These reference-aided architectures are fundamentally consistent with bistatic radar principles and clearly demonstrate the potential of Wi-Fi signals for radar-like sensing. However, due to their reliance on cable routing or dedicated reference antenna placement, most prior Wi-Fi radar functionalities cannot be integrated into typically deployed Wi-Fi systems.
These observations reveal a fundamental gap between existing Wi-Fi sensing and Wi-Fi radar approaches. While reference-aided Wi-Fi radar enables physically meaningful delay and Doppler analysis, such architectures cannot be integrated into typical Wi-Fi systems. Conversely, Wi-Fi sensing approaches that rely, for example, on CFR magnitude alone do not require any modification to existing Wi-Fi systems, but lack direct access to physically interpretable propagation parameters.
This raises a key question: Can bistatic Wi-Fi radar functionality be realized using only unmodified COTS Wi-Fi devices, without wired references or dedicated reference antennas, while being integrated into typical communication-oriented Wi-Fi systems, and still enabling physically meaningful delay–Doppler estimation?
@@
  • Unlike most existing Wi-Fi sensing approaches that operate directly on the CFR, which is originally acquired for communication purposes, a CIR is constructed using CFR obtained from COTS IEEE 802.11ax devices with $160\text{\,}\mathrm{MHz}$ bandwidth [18]. This is enabled by carefully designed frequency-domain preprocessing of the raw CFR, and provides a reliable basis for subsequent phase-coherent delay-domain processing.
  • To the best of our knowledge, this is the first work that realizes reference-aided bistatic Wi-Fi radar that can be integrated into typically deployed Wi-Fi systems, using only unmodified COTS Wi-Fi devices, without relying on wired references or dedicated reference antennas. To this end, we propose an over-the-air (OTA) line-of-sight (LoS)-path referencing scheme (LoSRef) that exploits the temporal stability of the transmitter–receiver (Tx–Rx) LoS path to provide a common delay and phase reference in practical Wi-Fi systems, as illustrated in Fig. 1(c), thereby enabling phase-coherent CIR acquisition.
  • The proposed LoSRef-based framework requires only a single antenna port at the Rx, whereas conventional reference-aided bistatic Wi-Fi radar functionalities typically rely on multi-port Rxs or dedicated reference channels. This further simplifies hardware requirements and facilitates integration into typical Wi-Fi systems using unmodified COTS Wi-Fi devices.
@@
This paper is organized as follows. Section II reviews related work on Doppler-based Wi-Fi sensing and reference-aided Wi-Fi radar. Section III discusses the reference ambiguity due to independent clocks in practical Wi-Fi systems. Section IV presents the proposed phase-coherent CIR construction framework, including delay and phase referencing based on a stable LoS-path reference. Section V demonstrates bistatic range estimation through human gait and respiration experiments. Section VI presents delay–Doppler response estimation results for moving targets. Section VII provides additional discussion, followed by Section VIII, which concludes the paper.
@@ II-A Doppler-Based Wi-Fi Sensing
TABLE I: Positioning of This Work with Respect to Existing Studies Including ISAC-Enabling Capability
Study Tx Rx Reference Signal (Fig. 1) ISAC-enabling Standard
Typical Wi-Fi sensing [24, 23, 25] COTS COTS $\times$ IEEE 802.11n
[10, 12, 13] COTS Instrument Rx (a) Wired reference $\times$ IEEE 802.11g
[9, 11] COTS Instrument Rx (b) Reference antenna $\times$ IEEE 802.11g
[14, 16] COTS COTS (a) Wired reference $\times$ IEEE 802.11n
[15] COTS COTS (b) Reference antenna $\times$ IEEE 802.11n
[17] COTS COTS (b) Reference antenna $\times$ IEEE 802.11ax
@@ II-B Reference-Aided Wi-Fi Radar
Table I summarizes the positioning of this work with respect to representative Wi-Fi sensing and reference-aided Wi-Fi radar studies, highlighting differences in Tx and Rx types, reference signals, ISAC-enabling capability, and supported standards. Here, ISAC-enabling indicates whether such functionality can be realized within typically deployed Wi-Fi systems without additional reference hardware. In [9, 11], a bistatic radar architecture is adopted, where a COTS Wi-Fi Tx and a measurement-grade Rx are used. The Rx employs separate reference and surveillance antennas, as illustrated in Fig. 1(b), with the antenna configuration designed such that target-induced variations are primarily captured in the Tx–surveillance path. The target-induced Doppler frequency is then extracted by correlating the reference and surveillance signals.
@@
In contrast to existing reference-aided Wi-Fi radar systems, which rely on wired references or dedicated reference antennas, the LoSRef exploits the OTA Tx–Rx LoS path. As a result, phase-coherent bistatic radar operation is achieved using only a single-antenna, single-port Rx, without requiring multi-port Rx architectures, thereby enabling integration into typical Wi-Fi systems and facilitating ISAC operation.
III CIR Representation and Reference Ambiguity Due to Free-Running Clocks
This section briefly reviews the CIR representation and highlights the reference ambiguity arising from free-running clocks in practical Wi-Fi systems, which fundamentally prevents phase-coherent analysis and physically meaningful time-domain processing.
@@ III-A CIR Representation and Doubly Selective Channels
III-B Reference Ambiguity Due to Free-Running Clocks
In Wi-Fi systems, where the Tx and Rx employ independent, free-running clocks, neither a common timing reference nor an absolute phase reference is available between them. As a result, the CIR observed at the Rx does not preserve the absolute delay reference to the transmission timing, nor does it retain a phase-coherent representation across successive observations. Accordingly, the observed CIR $h_{\mathrm{raw}}(\tau,t)$ is related to the underlying physical CIR $h(\tau,t)$ through an unknown delay offset and an unknown phase offset, and is given by
@@ III-B Reference Ambiguity Due to Free-Running Clocks
where $\alpha_{t}$ and $\beta_{t}$ denote the delay offset and phase offset, respectively, which may vary across observation times $t$ [21, 22]. These offsets originate from free-running clocks at the Tx and Rx, and prevent direct use of absolute delay and phase information, thereby precluding phase-coherent analysis of the doubly selective channel in practical Wi-Fi systems without a shared reference. Note that the delay offset $\alpha_{t}$ of the devices used in this study has been experimentally characterized in [17].
@@ IV-B LoSRef-Based Delay Calibration and Phase Alignment
Figure 3: LoSRef: OTA LoS-path referencing for delay calibration and phase alignment. The dominant LoS-path component is used as an OTA reference to compensate unknown offsets arising from free-running clocks at the Tx and Rx.
@@ IV-C Clutter Removal via LoSRef-based CIR Alignment
Note that, in conventional Wi-Fi sensing systems, direct processing of complex-valued CIRs is not physically meaningful due to the lack of a shared delay and phase reference between the Tx and Rx, which prevents meaningful temporal averaging. For example, Widar [25] extracts motion-related components by applying PCA to the CFR magnitude, followed by time–frequency spectral analysis. By contrast, the proposed LoSRef-based approach aligns CIRs on a common delay axis using a stable LoS-path reference, enabling phase-coherent temporal averaging and delay-domain clutter removal, which may be advantageous for slowly varying motions.
@@ VIII Conclusion
Through representative human gait and respiration experiments, we confirmed that phase-coherent bistatic radar functionality can be realized within typically deployed, communication-oriented Wi-Fi systems. Specifically, the proposed framework enables physically meaningful delay and Doppler estimates, including gait direction discrimination and respiration-induced sub-wavelength displacement measurements.